Find the lengths of each side give the following right triangle:
Base=x
HYPOTENUSE= x+8
last side = 2x-8

Answer by rothauserc(4718)

Use the Pythagorean Theorem
:
x^2 + (2x-8)^2 = (x+8)^2
:
x^2 + 4x^2-32x+64 = x^2+16x+64
:
5x^2-32x+64 = x^2+16x+64
:
4x^2-48x = 0
:
divide both sides of = by 4
:
x^2 -12x = 0
:
x(x-12) = 0
:
x = 0 or x = 12
:
we reject x = 0, since that means the base is 0 length and other side is negative length
:
*********************************************
base = 12, hypotenuse = 20, other side = 16
:
check answer
:
12^2 + 16^2 = 20^2
:
144 + 256 = 400
:
400 = 400
:
answer checks
***********************************************
: